Y hat = 20 - 2x, if we increase the value of x by 10, what kind of increase will we see in y?
Answer #1

Answer:

Given that:

Y hat = 20 - 2x, if we increase the value of x by 10, what kind of increase will we see in y?

Y hat = 20 - 2x, if we increase the value of x by 10

= 20 - 2x

if x =10 then

= 20 - 2(10)

= 20 - 20

= 0